Find the P - value for the test statistic ?=2.74z=2.74 for the following null and alternative hypotheses:
?0H0: The population mean is 5.
??Ha: The population mean is less than 5.
The P - value is
(b) Find the P - value for the test statistic ?=2.74z=2.74 for the following null and alternative hypotheses:
?0H0: The population mean is 5.
??Ha: The population mean is not equal to 5.
The P - value is

These are both one question for me!! Ive tried this one so many times help!!
Answer #1

solution

(A)this is left tailed test

P(z < 2.74 ) = 0.9969

P-value = 0.9969 using z table

(B)this is two tailed test

P(z > 2.74) = 1 - P(z < 2.74 ) = 1-0.9969=0.0031*2=0.0062

P-value = 0.0062 using z table